OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1). You must show all work.

OpenStudy (anonymous):

The formula for finding the equation of line when two points are given is (y-y1)=(x-x1)(y2-y1)/(x2-x1) Now y+6=(x-3)(-1+6)/(-2-3) y+6=(x-3)(-1) y+6=-x+3 x+y+3=0

OpenStudy (anonymous):

Nice job Atchyut!